OpenStudy (anonymous):

Help please!!! Simplify complex fraction ((y/y-3)+(y/1))/ (y-3/3

OpenStudy (anonymous):

multiply the numerator by (3/(y-3))

OpenStudy (anonymous):

I know the steps but I just don't understand HOW to do it. Im really confused

OpenStudy (anonymous):

\[\frac{ y }{ y-3 }+\frac{ y }{ 1 }\div \frac{ y-3 }{ 3 }\] is that the same?

OpenStudy (anonymous):

yes that's it. I have to simplify it

OpenStudy (anonymous):

i would let @tomo handle this, otherwise i would check back in

OpenStudy (anonymous):

okay, thanks

OpenStudy (anonymous):

ok, so we have ((y/y-3)+(y/1))/ (y-3/3) we can multiply the numerator by the inverse of the denominator. [y/(y-3) + (y/1)] * [3/(y-3)] [3y/(y-3)^2 + 3/(y-3)] ok so now you can factor out a 3/(y-3) [3/(y-3)] [(y/(y-3))+1]

OpenStudy (anonymous):

Thanks so much!!

OpenStudy (anonymous):

yw